OpenStudy (anonymous):

PLEASE HELP!!!! A candle is 17 in. tall after burning for 3 hours. After 5 hours, it is 15 in. tall. Write a linear equation to model the relationship between height h of the candle and time t

OpenStudy (anonymous):

A linear equation has the form y=mx + b where y is the dependant variable, x is the independent variable, m is the slope, and b is the y-intercept Let's change the variables to suit our situation. h = mt + b where h is the height and t is the time. Let's look at our two points given. I'll write it like (t,h) (3hrs, 17in) (5hr, 15in) Now let's plug those equations back into the h = mt + b \[17 = m*3 + b\] \[15 = m*5 + b\] Now we have a system of two equations and two unknowns. Solve for m and b. Once you get m, the slope, and b, the h-intercept, you can plug it back into the linear equation. h = mt + b. Voila, there's your equation

OpenStudy (anonymous):

I'm still confused. I see what you did, but is h = mt+b the equation?